OpenStudy (anonymous):

Please help! Find f(-x)-f(x) for the following function: f(x)=x^3+x-9

OpenStudy (mertsj):

What is f(-x)?

OpenStudy (anonymous):

The question didn't say and so I'm stuck.

OpenStudy (anonymous):

\(f(3)=3^2+3-9=27+3-9=21\) \(f(-3)=(-3)^2+(-3)-9=-27-3-9=-39\) \(-39-21=-60\) now try it with just \(x\) and \(-x\)

OpenStudy (anonymous):

\(f(-x)\) is confusing, but it means replace \(x\) everywhere by \(-x\)

OpenStudy (anonymous):

so first step is to write \(f(-x)= (-x)^3+(-x)-9\) and then simplify a big take that answer, and subtract \(x^3+x-9\) from it make sure to use parentheses

OpenStudy (anonymous):

So it would be -x-x-9-(x^2+x-9)?

OpenStudy (anonymous):

i assume you mean \[-x^3-x-9-(x^3+x-9)\]

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

ok good, now remove the parentheses by distributing the minus sign, then combine like terms

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

yw hope it is clear